OpenStudy (anonymous):

I need to know how to write a rational function by only knowing the vertical and horizontal asymptotes.

OpenStudy (anonymous):

do you have x and y?

OpenStudy (anonymous):

yes. All i have is Vertical asymptote at x = 5 and a horizontal asymptote at y = 1/2

OpenStudy (mertsj):

If there is a vertical asymptote at x = 5, then x-5 is a factor of the denominator. If there is a horizontal asymptote at y = 1/2, that tells you that the degrees of the numerator and denominator are the same and the ratio of the coefficients of the leading terms is 1/2

OpenStudy (anonymous):

ok now what?

OpenStudy (mertsj):

You're done

OpenStudy (anonymous):

but all i know is the denominator

OpenStudy (anonymous):

even read his comment ;)

OpenStudy (anonymous):

ya im confused

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

you should practice more read carefully what he write and practice the best thing is like if really you dont understand come again :)

OpenStudy (anonymous):

? i under stand what he is saying about the vertical asymptote but i don't understand the horizontal asymptote.

OpenStudy (mertsj):

Do you know what is meant by the degree of the numerator and denominator?

OpenStudy (anonymous):

yes x or x^2 or x^3 etc

OpenStudy (mertsj):

So, if I have a numerator of x^3-5x^2+6x-9, what is the degree of the numerator?

OpenStudy (anonymous):

x^3

OpenStudy (mertsj):

The degree is 3.

OpenStudy (anonymous):

ya

OpenStudy (mertsj):

So earlier I told you that if the degrees of the numerator and denominator are the same, then the horizontal asymptote is the ratio of the coefficients. Do you know what coefficients are?

OpenStudy (anonymous):

the number before the variables?

OpenStudy (mertsj):

And it is the ratio of the coefficients of the leading terms and the leading terms are the ones that determine the degree.

OpenStudy (mertsj):

So...an example

OpenStudy (mertsj):

\[f(x)=\frac{3x+2}{x-2}\]

OpenStudy (mertsj):

Do you see that the degrees of the numerator and denominator are equal?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

Do you see that the leading term of the numerator is 3x and of the denominator is x?

OpenStudy (anonymous):

yes so it would be 3/1?

OpenStudy (mertsj):

So the horizontal asymptote is y = 3. Yes. and there is a vertical asymptote at x=2

OpenStudy (mertsj):

Can you do your problem now?

OpenStudy (anonymous):

so would \[\frac{ x + 10 }{ 2x - 10 }\] work??

OpenStudy (mertsj):

That will work. Good for you.

OpenStudy (anonymous):

Thanks!

OpenStudy (mertsj):

yw